### v1

@@ 6 Conclusions
In conclusion, we introduce contrastive adversarial training as a novel approach to evaluating the robustness of LDM customization against protective perturbations. We first reveal that the primary reason adversarial examples serve as effective protective perturbations in LDM customization is the distortion of their latent representation. We then propose CAT, which employs adapters in adversarial training by leveraging contrastive adversarial loss. Extensive experiments demonstrate that CAT significantly decreases the effectiveness of existing protective perturbations. We believe this work can remind the community to reconsider the robustness of current protective perturbation methods.
Appendix A Additional Experimental Results
A.1 Qualitative Results for Object-Driven Image Synthesis in DreamBooth (CelebA-HQ)
A.2 Customization for Object-driven Image Synthesis in LoRA
Table 3: Quantitative results for object-driven image synthesis using CAT methods customized in DreamBooth. The table presents performance metrics FSS and FQS for the CelebA-HQ and VGGFace2 datasets across different protective perturbation settings. The optimal values among different CAT settings for each metric are bolded.
@@ A.2 Customization for Object-driven Image Synthesis in LoRA
sds+ 0.464 0.497 0.559 0.285 0.381 0.328 0.525 0.389 0.393 0.541 0.443 0.413 0.406 0.399 0.505 0.392
sds- 0.476 0.625 0.586 0.535 0.297 0.583 0.665 0.490 0.353 0.620 0.591 0.480 0.368 0.633 0.695 0.416
sdsT5 0.203 0.572 0.486 0.032 0.217 0.488 0.574 0.137 0.112 0.549 0.544 0.052 0.248 0.578 0.716 0.217
glaze2 0.355 0.669 0.575 0.479 0.474 0.714 0.752 0.617 0.330 0.602 0.611 0.472 0.576 0.673 0.755 0.586
antidb 0.539 0.596 0.616 0.465 0.497 0.458 0.541 0.411 0.460 0.583 0.586 0.439 0.416 0.569 0.725 0.494
@@
Appendix B Details of Customization Settings

### v2

@@ 6 Conclusions
In this work, we introduce Contrastive Adversarial Training (CAT) as a novel approach to evaluating the robustness of LDM customization against protective perturbations. We identify that the primary factor contributing to the effectiveness of adversarial examples as protective perturbations lies in the distortion of their latent representations. We then propose CAT, which employs lightweight adapters trained using the contrastive adversarial loss to realign the latent representations. Extensive experiments demonstrate that CAT significantly reduces the effectiveness of a range of existing protective perturbation methods in both object-driven synthesis and style mimicry customization tasks. We believe this work can encourage the community to reconsider the robustness of existing protective perturbations and inspire the development of more resilient protections in the future.
Acknowledgements
This work was supported by Hong Kong RGC Research Impact Fund with No.R1012-21 and General Research Fund with CityU No.11211422. We sincerely thank Linshan Hou from Harbin Institute of Technology, Shenzhen for the valuable comments and insightful suggestions.
Impact Statement
This work evaluates the robustness limitations of existing protective perturbation methods against unauthorized data usage in diffusion-based generative models. Our goal is to highlight key vulnerabilities and provide insights toward the design of more resilient protection mechanisms. In particular, our findings suggest that effective protection may benefit from incorporating defense strategies that explicitly target the diffusion process, especially within end-to-end optimization frameworks. In contrast, relying solely on latent autoencoders may be less robust, as they are more easily compromised by adaptive attacks. This perspective points to a possible direction for developing more robust protective perturbations that better safeguard the intellectual property rights of data owners. All visual content used in this study was either publicly licensed or provided with explicit permission.
Appendix A Visualization of Latent Representations using UMAP
Figure 8: Visualization of latent representations in UMAP of clean sample $z_{c}$, noisy sample $z_{r}$, protected sample $z_{a}$, and the latent representation of protected sample $z_{a}^{(\text{cat})}$ encoded by the latent encoder with CAT adapter. Protected samples are generated using nine protective perturbation methods, with four clean images from the VGGFace2 dataset as the base. The perturbation budget for protected and noisy samples is fixed at $\delta=16/255$. Each number labels the representations obtained from the same image.
Appendix B Details of Customization Settings
Appendix C Additional Experimental Results
C.1 Additional Quantitative Results for Object-Driven Image Synthesis in LoRA
Table 5: Quantitative results for object-driven image synthesis using CAT methods customized in LoRA. The table presents performance metrics FSS and FQS for the CelebA-HQ and VGGFace2 datasets across different protective perturbation settings. The optimal values among different CAT settings for each metric are bolded.
@@ C.1 Additional Quantitative Results for Object-Driven Image Synthesis in LoRA
SDS(+) 0.464 0.497 0.559 0.285 0.381 0.328 0.525 0.389 0.393 0.541 0.443 0.413 0.406 0.399 0.505 0.392
SDS(-) 0.476 0.625 0.586 0.535 0.297 0.583 0.665 0.490 0.353 0.620 0.591 0.480 0.368 0.633 0.695 0.416
SDST 0.203 0.572 0.486 0.032 0.217 0.488 0.574 0.137 0.112 0.549 0.544 0.052 0.248 0.578 0.716 0.217
Glaze 0.355 0.669 0.575 0.479 0.474 0.714 0.752 0.617 0.330 0.602 0.611 0.472 0.576 0.673 0.755 0.586
Anti-DB 0.539 0.596 0.616 0.465 0.497 0.458 0.541 0.411 0.460 0.583 0.586 0.439 0.416 0.569 0.725 0.494
@@
C.2 Additional Quantitative Results for Object-Driven Image Synthesis in DreamBooth
Table 6: Quantitative results for object-driven image synthesis using CAT methods customized in DreamBooth, compared to the baseline across CelebA-HQ and VGGFace2 datasets. The table reports FID scores, where lower values indicate better alignment with the real distribution. Optimal values among CAT settings are bolded for each dataset and method.
FID $\downarrow$ CelebA-HQ VGGFace2
Method Baseline CAT-both CAT-en Baseline CAT-both CAT-en
AdvDM(+) 340.0 264.9 223.7 435.2 274.9 249.0
AdvDM(-) 134.3 104.0 102.0 203.9 189.4 188.6
Mist 263.6 133.6 136.1 359.6 187.8 198.9
SDS(+) 327.2 277.4 247.6 363.9 295.4 255.7
SDS(-) 125.4 103.4 110.2 208.9 183.3 185.6
SDST 223.0 133.3 133.5 335.8 195.3 200.4
Glaze 196.7 100.1 90.4 228.0 160.9 191.3
Anti-DB 180.4 131.4 106.4 320.5 202.1 190.8
MetaCloak 175.0 179.6 171.6 316.3 200.4 170.9
C.3 Additional Qualitative Results for Object-Driven Image Synthesis in DreamBooth